# Contest for system observability as an infinitely repeated game

### v2

@@ 4.1 Game based on one-step optimality
(Sufficiency) Because $\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega$ $(U_{a}^{*},H)=\operatorname{dim\,Ker}\Omega(U_{a}^{*},H^{*})$, $H^{*}\in BR1^{d}$. Since $\operatorname{dim\,}\mathcal{V}^{*}(H^{*})$ $=\operatorname{dim\,Ker}\Omega(U_{a}^{*},H^{*}),U_{a}^{*}$ is a friend matrix of $\mathcal{V}^{*}(H^{*})$ and $U_{a}^{*}\in BR1^{a}$. Thus $(U_{a}^{*},H^{*})$ is a Nash equilibrium. (Necessity) Given the symmetry in the argument, the necessity part is omitted.
@@
Theorem 4.5 is effective in determining when the Nash equilibrium is reached in dynamic game process, as it only requires verifying if $\,\min_{{H}\in\mathbb{R}^{m\times n}}\!\operatorname{dim\,Ker}\Omega\,(U_{a}^{*},\!H)$ $=\!\operatorname{dim}\mathcal{V}^{*}(H^{*})$, i.e., whether the value function after the defender chooses $H^{*}$ is equal to $\operatorname{dim}\mathcal{V}^{*}(H^{*})$. If this condition is satisfied, since $\,\min_{{H}\in\mathbb{R}^{m\times n}}\!\operatorname{dim\,Ker}\Omega\,(U_{a}^{*},\!H)=\operatorname{dim\,Ker}\Omega\,(U_{a}^{*},H^{*})$ always holds when both players consider one-step optimality, the condition in Theorem 4.5 is satisfied and the game reaches Nash equilibrium.
Figure 3: An example of the evolution of best response sets and actions. Coordinate axis represents game epochs. Different shapes above the axis symbolize different best response sets per epoch. Differently colored points represent different actions; those below the axis represent the actions chosen in each epoch.
@@
Next we want to analyze the game result in dynamic game process. Since the best response sets for both players are not single-valued maps, we need to add assumptions on the choices for both players:
@@
To further illustrate the above assumptions, Fig. 3 represents an example of the evolution of the best response sets and actions for one player, who updates its action every two epochs. In epoch 1, the player chooses an action (the red point) in the best response set (the square pattern). Then in epoch $k+2$, the player has the same best response set as that in epoch 1 and chooses the same action as that in epoch 1 according to Assumption 2. Afterwards, in epoch $m+2$, though the player has the same best response set as that in epoch 1, it keeps the action unchanged (the blue point), because the last action (the blue point) also belongs to the best response set in this epoch and Assumption 1 is prioritized over Assumption 2.
@@
Assumptions 1 and 2 frequently manifest in real-world applications. Assumption 1 can save energy used in changing actions, while Assumption 2 fixes the rule to select a solution in a multi-valued map (such as choosing the one with the smallest norm length).
Based on these assumptions, we can easily get the following three corollaries which represent different game results.
Corollary 4.8.
For infinitely repeated game considering one-step optimality, the value function is locked, i.e., $\exists l\in\mathbb{N}$ and $\gamma\in\mathbb{R}$, in epoch $i\geq l,\operatorname{dim\,Ker}\Omega(U_{ai},H_{i})=\gamma$, if and only if players’ strategy profile in epoch $l$ $(U_{al},H_{l})$ is a Nash equilibrium.
@@
This corollary is nontrivial when the best response sets are not single-valued maps, because different choices in the best response set affect the upper bound of the subsequent action of another player. Therefore we need the assumptions 1-2.
Corollary 4.10.
If in epoch $l$, players’ strategy profile $(U_{al},H_{l})$ is not a Nash equilibrium, the value function oscillates and the amplitude equals to $\operatorname{dim\,}\mathcal{V}^{*}(H_{l})-\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega\,(U_{al},H).$
Corollary 4.11.
If the attacker or the defender chooses the same action in epoch $i$ and epoch $j,i<j$ with $(j-i)$ even, then the value function and players’ actions will evolve in a circle and $(j-i)$ is a multiple of the number of epochs in a minimal circle.
Proof 4.12.
Assume $U_{ai}=U_{aj}$. In epoch $j$, $(N+B_{a}^{\prime}U_{aj})$ returns to the same value as $(N+B_{a}^{\prime}U_{ai})$. Thus in epoch $(j+1)$, the $BR1^{d}$ is the same as $BR1^{d}$ in epoch $(i+1)$. According to Assumption 2, the defender will choose the same action in epoch $(j+1)$ as in epoch $(i+1)$. (Here we do not need to consider the scenario where Assumption 1 is satisfied, because in this case, the game reaches Nash equilibrium, which is also a circle.) Similarly, at each subsequent epoch, the attacker (or defender) will take the same action as $(j-i)$ epochs before, which leads to the circle of the value function and players’ actions. Thus $(j-i)$ is a multiple of the number of epochs in a minimal circle. If $H_{i}=H_{j}$, the proof is similar.
4.2 Game based on two-step optimality
The best response sets considering one-step optimality are not single-valued maps. Different initial preferences for choices based on Assumption 2 can lead to different game outcomes, which is showed in Section 5 Illustrative examples. This is because on the one hand, different values of $H$ in the set $BR1^{d}$ have different values of $\operatorname{dim\,}\mathcal{V}^{*}$, which determines the upper bound of $U_{a}$ to maximize the value function in the next epoch. On the other hand, different values of $U_{a}$ in the set $BR1^{a}$ will influence the maximal geometric multiplicity of $(N+B_{a}^{\prime}U_{a})$, which influences the lower bound of $\min_{H}\operatorname{dim\,Ker}\Omega\,(U_{a},H)$ in the next epoch. Thus $\operatorname{dim\,}\mathcal{V}^{*}$ and the maximal geometric multiplicity of $(N+B_{a}^{\prime}U_{a})$ are taken into consideration in this section.
Formulate the problem based on two-step optimality, which means two players are smart enough to choose an action in $BR1^{a}$ or $BR1^{d}$ that gives the opponent the worst outcome in the next step. There are two steps to find the best response sets following two-step optimality: $(i)$ find $BR1^{a}$ or $BR1^{d}$; $(ii)$ in $BR1^{a}$ or $BR1^{d}$, find the set of actions that make the opponent get the worst outcome in the next epoch.
The best response set for the attacker built on two-step optimality is
$\begin{split}BR2^{a}&=\arg\max_{U_{a}\in BR1^{a}}\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega\,(U_{a},H)\\ &=\arg\max_{U_{a}\in BR1^{a}}[\text{MGM}(N+B_{a}^{\prime}U_{a})].\end{split}$ (30)
where $\text{MGM}(N+B_{a}^{\prime}U_{a})$ is the maximum geometric multiplicity associated with the eigenvalues of $(N+B_{a}^{\prime}U_{a})$. This result can be directly derived from Corollary 3.3. Define $BR2X^{a}$ as

### v3

@@ 4.2 Game based on two-step optimality
$\forall\,H_{0}\in\mathbb{R}^{m\times n}$, the maximal controlled invariant subspace in $\operatorname{Ker}H_{0}$ is larger than $(N+B_{a}^{\prime}U_{a})$-invariant subspace in $\operatorname{Ker}H_{0}$, i.e., $\operatorname{dim\,}\mathcal{V}^{*}(H_{0})\geq\operatorname{dim\,Ker}\Omega(U_{a},H_{0}).$ Thus $\forall H_{1}\in arg\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,}\mathcal{V}^{*}(H)$, $\operatorname{dim\,}\mathcal{V}^{*}(H_{1})\geq\operatorname{dim\,Ker}\Omega(U_{a},H_{1})$. And $\forall H_{2}\in\arg\min_{{H}\in\mathbb{R}^{m\times n}}\,$ $\operatorname{dim\,Ker}$ $\Omega(U_{a},H)$, there is $\operatorname{dim\,Ker}\Omega(U_{a},H_{1})\geq\operatorname{dim\,Ker}\Omega(U_{a},H_{2})$. Thus $\operatorname{dim\,}\mathcal{V}^{*}(H_{1})\!\geq\!\operatorname{dim\,Ker}\Omega(U_{a},H_{2})$, which completes the proof.
Building on this lemma, we demonstrate that the Nash-equilibrium criterion stated in Theorem 4.2 also holds when both players adopt a two-step optimality framework. Although the equilibrium condition is identical for the one-step and two-step settings, the underlying mathematical reasoning are distinct.
Theorem 4.7 (Two-step optimality NE criterion).
The strategy profile $(U_{a}^{*},H^{*})$ is a Nash equilibrium when two players consider two-step optimality if and only if,
$\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega\,(U_{a}^{*},H)=\operatorname{dim\,}\mathcal{V}^{*}(H^{*}).$ (26)
Proof 4.8.
(Sufficiency) According to Theorem 4.2, $H^{*}\in BR1^{d}(U_{a}^{*})$, $U_{a}^{*}\in BR1^{a}(H^{*})$ and
$\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega\,(U_{a}^{*},H)=\operatorname{dim\,}\mathcal{V}^{*}(H^{*})=\operatorname{dim\,Ker}\Omega\,(U_{a}^{*},H^{*})\overset{\triangle}{=}\gamma.$ (27)
We need to further prove $H^{*}\in BR2^{d}(U_{d}^{*})$ and $U_{a}^{*}\in BR2^{a}(H^{*})$. Firstly, according to Lemma 4.5 and (27), $\forall H^{\prime}\in BR1^{d}(U_{a}^{*})$ we have
$\operatorname{dim\,}\mathcal{V}^{*}(H^{\prime})\geq\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,}\mathcal{V}^{*}(H)\geq\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega(U_{a}^{*},H)=\gamma.$ (28)
By (27), $\operatorname{dim\,}\mathcal{V}^{*}(H^{*})=\gamma$, which reaches the lower bound of the $\operatorname{dim\,}\mathcal{V}^{*}(H\in BR1^{d}(U_{a}^{*}))$. Thus $H^{*}\in arg\min_{H\in BR1^{d}(U_{a}^{*})}\operatorname{dim\,}\mathcal{V}^{*}(H)=BR2^{d}(U_{a}^{*})$. Secondly, by Lemma 4.5 and (27), $\forall U_{a}^{\prime}\in BR1^{a}(H^{*})$, there is
$\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega(U_{a}^{\prime},H)\leq\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,}\mathcal{V}^{*}(H)\leq\operatorname{dim\,}\mathcal{V}^{*}(H^{*})=\gamma.$ (29)
By (27), $\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega(U_{a}^{*},H)=\gamma$, which reaches the upper bound of $\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}$ $\Omega(U_{a}^{\prime},H)$. Thus $U_{a}^{*}\in\arg\max_{U_{a}\in BR1^{a}(H)}\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega\,(U_{a},H)=BR2^{a}(H^{*})$. Thus $(U_{a}^{*},H^{*})$ is the Nash equilibrium of the two-step optimal game.
(Necessity) Since $BR2^{d}\subseteq BR1^{d}$ and $BR2^{a}\subseteq BR1^{a}$, we have $H^{*}\in BR1^{d}(U_{a}^{*})$, $U_{a}^{*}\in BR1^{a}(H^{*})$. The rest proof is the same as Theorem 4.2.
Theorem 4.7 reveals that, although the two-step best response sets are a subset of the one-step best response sets and the strategy sets of both players differ, the resulting equilibrium points coincide.
@@
In this paper, the analysis is limited to one-step and two-step optimality. Extending the horizon to three or more steps introduces no additional strategic content, because the same player acts in both epochs 1 and 3, making the strategic situation in epoch 3 identical to that in epoch 1. Hence, higher‐order optimality criteria can be omitted without loss of generality, and the extended-horizon study in Section 4.2 (two-step optimality) exhausts all non-trivial multi-step cases.
4.3 Game outcome analysis
This subsection gives analysis for game results in dynamic game process that are consistent with both one-step and two-step optimality game. First, we examine a degenerate scenario in which the game possesses an equilibrium where the defender holds an absolute advantage. The following lemma states sufficient conditions under which this situation occurs.
Theorem 4.10 (Defender-dominated NE).
Let $N\in\mathbb{R}^{n\times n}$ and $B_{a}^{\prime}\in\mathbb{R}^{n\times k}$. Assume there exists $H^{*}\in\mathbb{R}^{m\times n}$ such that either of the following two cases holds:
Case 1: $(n-m)\geq k.\text{ If }\operatorname{Im}B_{a}^{\prime}\subseteq\operatorname{Ker}H^{*}$ and there is no nontrivial $N$-invariant subspace contained in $\operatorname{Ker}H^{*}$;
Case 2: $(n-m)<k$. If $\operatorname{Ker}H^{*}\subseteq\operatorname{Im}B_{a}^{\prime}$ and any vector in $\operatorname{Ker}H^{*}$ does not belong to $N$-invariant subspace contained in $\operatorname{Im}B_{a}^{\prime}$. Then
$\operatorname{dim}\mathcal{V}^{*}(H^{*})=0,$ (30)
and, for every $U_{a}\in\mathbb{R}^{k\times n}$, the strategy profile $(U_{a},H^{*})$ is a Nash equilibrium.
Proof 4.11.
[…]
Furthermore, when dim $\mathcal{V}^{*}(H^{*})=0$, $\forall U_{a}\in\mathbb{R}^{k\times n}$, $\operatorname{dim\,Ker}\Omega$ $(U_{a},H^{*})=\operatorname{dim\,}\mathcal{V}^{*}(H^{*})=0$. Thus $H^{*}\in\arg\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega$ $\,(U_{a},H)$. Plus, there is $\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega\,(U_{a},H)=\operatorname{dim\,}\mathcal{V}^{*}$ $(H^{*})=\operatorname{dim\,Ker}\Omega(U_{a},H^{*})$. According to Theorem 4.2 and Theorem 4.7, $(U_{a},H^{*}),\forall U_{a}\in\mathbb{R}^{k\times n}$ is an equilibrium whenever two players consider one-step or two-step optimality, which completes the proof.
When the conditions in Theorem 4.10 hold, the defender can choose a matrix $H^{*}$ such that $\mathcal{V}^{*}(H^{*})=\{0\}$. In this defender-dominated situation the attacker can no longer influence the system’s observability, and the game settles at a trivial equilibrium completely controlled by the defender. To exclude this degenerate case, the remainder of the paper focuses on the non-trivial regime $\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,}\mathcal{V}^{*}>0$. The following theorem provides a convenient sufficient condition under which this inequality is guaranteed.
Corollary 4.12 (Non-degenerate condition).
Consider system (7). Assume that $B_{a}^{\prime}\in\mathbb{R}^{n\times k}$ and $H\in\mathbb{R}^{m\times n}$ are both full-column-rank. If
$\max\bigl{\{}k,\operatorname{MGM}(N_{\overline{r}})\bigr{\}}>m,$ (31)
where $\operatorname{MGM}(N_{\overline{r}})$ denotes the maximum geometric multiplicity among all eigenvalues of the uncontrollable block $N_{\overline{r}}$ of the state matrix $N$, then
$\min_{H\in\mathbb{R}^{m\times n}}\dim\mathcal{V}^{*}(H)>0.$ (32)
Proof 4.13.
Split the system into controllable and uncontrollable parts: $N=\text{diag}(N_{r},N_{\overline{r}})$, $B_{a}^{\prime\!\top}=(B_{ar}^{\prime\!\top},0)$, and $H=(H_{r},H_{\overline{r}})$. Hence $\mathcal{V}^{*}=\mathcal{V}_{r}^{*}\oplus\mathcal{V}_{\overline{r}}^{*}$ with $\mathcal{V}_{r}^{*},\mathcal{V}_{\overline{r}}^{*}$ determined by the subsystems $(N_{r},B_{ar}^{\prime},H_{r})$ and $(N_{\overline{r}},0,H_{\overline{r}})$, respectively. Consider the following two cases. Case 1: If $k\geq\text{MGM}({N}_{\overline{r}})$, there is $k>m$ by (31). For controllable part $(N_{r},B^{\prime}_{ar},H_{r})$, the geometric multiplicity of $(N_{r}+B^{\prime}_{ar}U_{ar})^{\prime}s$ eigenvalue $\lambda$ is $\{n-\text{rank}[(N_{r}+B^{\prime}_{ar}U_{ar})-\lambda I]\}.$ According to PBH controllability criterion, for the controllable part we have $\forall\lambda\in\sigma(N_{r}+B^{\prime}_{ar}U_{ar})$, rank$[(N_{r}\!+\!B^{\prime}_{ar}U_{ar}\!-\!\lambda I)\,\,\,B^{\prime}_{ar}]=n$. Thus rank$[(N_{r}\!+\!B^{\prime}_{ar}U_{ar}\!-\!\lambda I)]\geq(n-k)$ and $\text{MGM}(N_{r}\!+\!B^{\prime}_{ar}U_{ar})=k.$ If $k>m$, i.e., $\text{MGM}(N_{r}\!+\!B^{\prime}_{ar}U_{ar})>m$, according to Proposition 3.3, =$\mathcal{I}=\{i\mid\alpha_{i}>m\}\neq\emptyset,\underset{{H_{r}}\in\mathbb{R}^{m\times n}}{\min}\operatorname{dim\,Ker}\Omega(U_{ar},H_{r})=n-\underset{{q}\in\mathcal{I}}{\min}(\sum_{p=1}^{m}r_{qj_{p}})>0.$ According to Lemma 4.5, $\min_{H_{r}\in\mathbb{R}^{m\times n}}\operatorname{dim\,}\mathcal{V}_{r}^{*}\geq\min_{H_{r}\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega(U_{ar},H_{r})>0$. Thus $\min_{{H}_{r}}\operatorname{dim\,}$ $\mathcal{V}_{r}^{*}>0$. Case 2: If $\text{MGM}({N}_{\overline{r}})>k$, there is $\text{MGM}({N}_{\overline{r}})>m$ by (31). Since $\mathcal{V}^{*}_{\overline{r}}=\operatorname{Ker}\Omega(N_{\overline{r}},0,H_{\overline{r}})$ for uncontrollable subspace, if $\text{MGM}({N}_{\overline{r}})>m$, extending Lemma 3.2 to the dual system, we have $\min_{{H}_{\overline{\mathrm{r}}}}\operatorname{dim\,Ker}\Omega>0$. Thus $\min_{{H}_{\overline{\mathrm{r}}}}\operatorname{dim\,}\mathcal{V}_{\overline{r}}^{*}>0$. To conclude, in either case, there is $\operatorname{dim\,}\mathcal{V}^{*}=\operatorname{dim\,}\mathcal{V}_{r}^{*}+\operatorname{dim\,}\mathcal{V}_{\overline{r}}^{*}>0,$ which completes the proof.
@@ 4.3 Game outcome analysis
Next, under the non-degenerate condition, we analyze the outcome of the game. Since the best response sets for both players are not single-valued mappings, we impose the following assumptions on strategy selection of both players:
@@
To further illustrate the above assumptions, Figure 4 represents an example of the evolution of the best response sets and actions for one player, who updates its action every two epochs.
Figure 4  An illustration of how best response sets and chosen actions evolve over game epochs: shapes above the axis denote the best response set of each epoch, while colored points below indicate the action selected at that epoch.
In epoch 1, the player chooses an action (the red point) in the best response set (the square pattern). Then in epoch $k+2$, the player has the same best response set as that in epoch 1 and chooses the same action as that in epoch 1 according to Assumption 2. Afterwards, in epoch $m+2$, though the player has the same best response set as that in epoch 1, it keeps the action unchanged (the blue point), because the last action (the blue point) also belongs to the best response set in this epoch and Assumption 1 is prioritized over Assumption 2.
@@
Assumptions 1 and 2 frequently manifest in real-world applications. Assumption 1 can save energy used in changing actions, while Assumption 2 fixes the rule to select a solution in a multi-valued map (such as choosing the one with the smallest norm length). Based on these assumptions, we can easily get the following theorem which represents two special game results.
Theorem 4.14 (Game results analysis).
Under Assumptions 12, there are two possible outcomes in the infinitely repeated game when considering either one-step or two-step optimality:
  1. (i)
    Lock mode. $\exists\,l,\gamma\in\mathbb{R}$, $\forall i\geq l$ such that $U_{a,i}=U_{a,l}$, $H_{i}=H_{l}$ and $\dim\operatorname{Ker}\Omega(U_{a,i},H_{i})=dim\operatorname{Ker}\Omega(U_{a,l},H_{l})$ if and only if $(U_{a,l},H_{l})$ is a Nash equilibrium.
  2. (ii)
    Loop mode. Both the strategy profile $(U_{a},H)$ and the value function $\dim\operatorname{Ker}\Omega$ evolve on a finite cycle, if and only if either player repeats an action after an even number of epochs, i.e. $U_{a,i}=U_{a,j}$ or $H_{i}=H_{j}$, whose minimal period divides $(j-i)$.
Proof 4.15.
(i) (Sufficiency) Assume that the profile $(U_{a,l},H_{l})$ is a Nash equilibrium, i.e. $U_{a,l}\in BRi^{a}(H_{l}),\,H_{l}\in BRi^{d}(U_{a,l})$, $i=1\text{ or }2$. Because each player is already playing a best response, by Assumption 1, both players prefer to keep the action unchanged, and inductively in every epoch $i\geq l$. (Necessity) Conversely, suppose $\forall i\geq l$ such that $U_{a,i}=U_{a,l}$, $H_{i}=H_{l}$. If $(U_{a,l},H_{l})$ were not a Nash equilibrium, at least one player would have a unilateral deviation in epoch $l+1$, contradicting that the strategy profile $\{U_{a,i},H_{i}\}$ remains unchanged. Hence $(U_{a,l},H_{l})$ must be an equilibrium.
(ii) (Sufficiency) Assume that the attacker repeats an action, i.e. $U_{a,i}=U_{a,j}$ with $j-i$ even. Then $N+B_{a}^{\prime}U_{a,j}=N+B_{a}^{\prime}U_{a,i}$, so the best response set of the defender in epoch $j+1$ is identical to that in epoch $i+1$. By Assumption 2, the defender therefore chooses the same action in epoch $j+1$ as in epoch $i+1$. Repeating the argument inductively, we find that each epoch reproduces the action taken $(j-i)$ periods earlier. Hence the entire strategy profile and the value function evolve on a finite cycle whose minimal period divides $j-i$. The proof is similar when the defender repeats an action, i.e., $H_{i}=H_{j}$. (Necessity) Since the strategy profile $(U_{a},H)$ evolves on a finite cycle, both players repeat the action after the period of the cycle. This completes the proof.
4.4 Equilibrium characterization
Although Theorems 4.2 and 4.7 provide a unified set of necessary and sufficient conditions for equilibrium, the criterion couples the two players’ strategies, which makes computing the equilibrium directly from the theorem impractical. Therefore, we derive the following necessary condition for the Nash equilibrium.
Theorem 4.16 (Necessary condition for NE).
For any Nash equilibrium $(U_{a}^{*},H^{*})$ when considering either one-step or two-step optimality, there must be
$H^{*}\in\arg\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,}\mathcal{V}^{*}(H),\,U_{a}^{*}\in\mathcal{F}(\mathcal{V}^{*}(H^{*})).$ (33)
Proof 4.17.
According to Theorem 4.2 and 4.7, for Nash equilibrium $(U_{a}^{*},H^{*})$, there is $\min_{{H}\in\mathbb{R}^{m\times n}}$ $\operatorname{dim\,Ker}\Omega\,(U_{a}^{*},\!H)=\operatorname{dim}\mathcal{V}^{*}(H^{*}).$ By further considering Lemma 4.5, we obtain
$\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,}\mathcal{V}^{*}(H)\geq\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega(U_{a}^{*},H)=\operatorname{dim}\mathcal{V}^{*}(H^{*}).$ (34)
Since $\operatorname{dim}\mathcal{V}^{*}(H^{*})\geq\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,}\mathcal{V}^{*}(H)$, the $\geq$’ symbol in (34) in fact holds with equality. Thus $H^{*}\in\arg\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,}\mathcal{V}^{*}(H)$. Plus, when considering either one-step or two-step optimality, there must be $U_{a}^{*}\in BR1_{a}(H^{*})=\mathcal{F}(\mathcal{V}^{*}(H^{*}))$. The proof is completed.
@@ 4.4 Equilibrium characterization
Since $\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,}\mathcal{V}^{*}(H)$ in (33) has no relevant to $U_{a}^{*}$, we can compute $H^{*}$ first and then choose $U_{a}^{*}$ which belongs to $\mathcal{F}(\mathcal{V}^{*}(H^{*}))$. Next, we test whether $(U_{a}^{*},H^{*})$ satisfies $\min_{H\in\mathbb{R}^{m\times n}}\dim\operatorname{Ker\,}\Omega(U_{a}^{*},H)=\dim\mathcal{V}^{*}(H^{*})$. If so, $(U_{a}^{*},H^{*})$ is a Nash equilibrium; otherwise, we select another pair $(U_{a}^{*},H^{*})$ that satisfies (33) and repeat the above procedure until an equilibrium is found.
In fact, sometimes Nash equilibrium is disadvantageous to the attacker because the value function remains at a low value that the attacker cannot change. In this case, the attacker can break the equilibrium by forsaking the current gain. Here follows a corollary.
Corollary 4.19 (Attacker-dominated non-equilibrium).
Assume $\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim}\mathcal{V}^{*}(H)>0$, which guarantees the existence of at least one attacker strategy $U_{a}^{*}$ such that